theyadoreshayyy:

If cos∠B = sin∠A and m∠A = 32°, what is the measure of ∠B? 32° 58° 90° 148°

theyadoreshayyy:

@tyrion

new2luv:

is there an image

theyadoreshayyy:

nope

Tyrion:

dats weird to have no image

Tyrion:

i would think u would need one

theyadoreshayyy:

idk

Tyrion:

do uk how to do this like at all

theyadoreshayyy:

no, i dont understand how to do it

tiahnadomino:

It is B. 58 because consider the right triangle ABC, with C as the right angle. The sides opposite those angles are a, b, c, where c is the hypotenuse. So if angle A is 32 then Angle B will be 90-32=58. For better understanding sin(A) = cos(90°-A) = cos(B) 90° -32° = B 58°

theyadoreshayyy:

thank yu

tiahnadomino:

you're welcome